# Approximate independence of permutation mixtures

### v1

@@ 1 Introduction
Compound decisions and empirical Bayes. In the compound decision problem [Rob51], the statistician observes independent data $X_{1},\dots,X_{n}$, where $X_{i}\sim P_{\theta_{i}}$ for $i=1,\dots,n$; the goal is to estimate the vector of parameters under a separabl loss $L(\theta,\widehat{\theta}):=\sum_{i=1}^{n}\ell(\theta_{i},\widehat{\theta}_{i})$. This setting is the starting point of the theory of empirical Bayes [Efr19], which views the parameters as having been drawn i.i.d. from a prior distribution, which is in turn estimated from the observations. An oracle version of this problem was suggested by Brown and Greenshtein [BG09] (see also [JZ09]), in which the statistician knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$, but not their correspondence to the observations $(X_{1},\dots,X_{n})$. Equivalently, the observations may be viewed as arising from the permutation mixture $\mathbb{P}_{n}$. As Greenshtein and Ritov [GR09] observe, however, most oracle inequalities in the empirical Bayes literature compare to an oracle that knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$ but is restricted to the use of “simple” decision rules, of the form $\hat{\theta}=(\Delta(X_{1}),\dots,\Delta(X_{n}))$ for some fixed univariate function $\Delta$. Such rules are in fact optimal when the observations arise from not from $\mathbb{P}_{n}$ but from its independent counterpart $\mathbb{Q}_{n}$. The empirical Bayes literature contains several ad hoc techniques for bounding the difference between these two approaches, going back to [HR55]. Our results can be used to give new, general risk bounds in this setting, see Corollary 1.7.
@@
Information geometry of high-dimensional mixture models. High-dimensional mixtures—such as the measures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ defined above—are ubiquitous models in theoretical statistics (see, e.g., [LNS99, CL11, IS12, WY16, JHW18]. The analysis of such measures is routine in the univariate case (or, more generally, for product measures), but obtaining sharp bounds is difficult for high-dimensional mixtures whose coordinates are not independent. The setting we study is one in which we can rigorously compare $\mathbb{P}_{n}$ to a simpler, independent counterpart $\mathbb{Q}_{n}$. This comparison may be directly relevant in analyzing mixtures with permutation invariance. Moreover, as we highlight in Section 1.2, our proofs are based on new techniques for bounding divergences between high-dimensional mixtures, which may be of broader applicability.
@@
    If $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$, then
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1.$ (5)
@@
Another direct corollary of Theorem 1.2 is the application of the continuity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ to permutation priors. Let $P_{\theta_{1}},\dots,P_{\theta_{n}}$ be fixed distributions over a space ${\mathcal{X}}$, and $\vartheta=(\vartheta_{1},\dots,\vartheta_{n})$ be a uniformly random permutation of $(\theta_{1},\dots,\theta_{n})$, i.e. $\vartheta_{i}=\theta_{\pi(i)}$ with $\pi\sim\text{\rm Unif}(S_{n})$. In addition, let $X=(X_{1},\dots,X_{n})$ be a random vector with independent coordinates $X_{i}\sim P_{\vartheta_{i}}$ conditioned on $\vartheta$. The next corollary states that, the mutual information $I(\vartheta;X)$ is close to the counterpart where the permutation prior of $\vartheta$ is replaced by an i.i.d. prior.
@@
Under the above setup, it holds that
@@
$\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\Delta_{H^{2}}(\{P_{\theta_{1}},\dots,P_{\theta_{n}}\}))\right\}.$
In words, Corollary 1.4 implies that the difference in the mutual information $I(\vartheta;X)$ under a permutation prior and an i.i.d. prior of $\vartheta$ is typically independent of $n$, i.e. these priors are “equally informative” for the channel output $X$ in the asymptotic sense. This implication turns out to find statistical applications for problems with permutation priors; we refer to Section 5.2 for such an example in the Gaussian sequence model.
@@
$\chi^{2}\left(P_{Y^{k}}\|Q_{Y^{k}}\right)\leq\frac{k(k-1)}{n(n-1)}\cdot\begin{cases}10\sum_{\ell=2}^{n}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell}&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})<\infty,\\ (e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1&\text{if }\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty.\end{cases}$
In the absence of a smoothing kernel, some well-known results due to Stam [Sta78], Diaconis–Freedman [DF80], and several recent works [GK21, JGK24] yield bounds of the form ${\sf KL}(P_{X^{k}}\|Q_{X^{k}})=O(\frac{k^{2}}{n})$, which can be improved to $O(\frac{k^{2}}{n(n+1-k)})$ when $|{\mathcal{X}}|=O(1)$. In particular, $P_{X^{k}}$ and $Q_{X^{k}}$ are, in general, statistically indistinguishable when $k=o(\sqrt{n})$, and that this requirement can be improved to $k=o(n)$ when ${\mathcal{X}}$ is small. [^1] Note that if no restriction is made on the size of ${\mathcal{X}}$, then the birthday paradox shows that the requirement $k=o(\sqrt{n})$ is tight. In comparison, Theorem 1.5 implies that as long as the exchangeable distribution can be written as a convolution with a kernel ${\mathsf{K}}$ satisfying $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$ (such as a Gaussian), the mixtures are statistically indistinguishable for any $k=o(n)$, without any restriction on $|{\mathcal{X}}|$. This observation suggests that the addition of noise in the form of the kernel ${\mathsf{K}}$ has a similar qualitative effect as a bound on the size of ${\mathcal{X}}$; we leave investigation of this interesting connection to future work.
@@
$\left(\mathbb{E}_{\mathbb{P}_{n}}[f]-\mathbb{E}_{\mathbb{Q}_{n}}[f]\right)^{2}\leq\mathbb{E}_{\overline{P}^{\otimes n}}\left[f^{2}\right]\cdot\frac{12{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{6{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n}.$
In other words, Theorem 1.6 shows that the squared TV distance between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of the order $O_{\mathcal{P}}(\frac{1}{n})$. In addition, the mean difference of any test function $f$ under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of order $O_{\mathcal{P}}(n^{-1/2})$ as long as the second moment of $f$ under the i.i.d. distribution $\overline{P}^{\otimes n}$ is bounded.
@@
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq M\sqrt{6n{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{6{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}.$
In other words, Corollary 1.7 shows that for a large class of models and losses, asymptotically these two types of decision rules achieve similar mean risks $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(\sqrt{n})$, strengthening a “folklore” result $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=o(n)$ in [HR55]. For the quadratic loss $L(\vartheta,\widehat{\vartheta})=\|\vartheta-\widehat{\vartheta}\|_{2}^{2}$, the risk difference can be improved to $O(1)$ via a different strategy, as shown for a class of models in [GR09]; we will discuss this improvement in Section 5.3.
@@ 1.2 Proof techniques
$|e_{\ell}(x)|\leq C\sqrt{\binom{n}{\ell}}\left(\sum_{i=1}^{n}x_{i}^{2}\right)^{\ell/2}.$ (8)
This bound may be contrasted with Maclaurin’s inequality [Mac30], a version of which implies the bound $|e_{\ell}(x)|\leq\binom{n}{\ell}\left(\sum_{i=1}^{n}x_{i}^{2}\right)^{\ell/2}$ for all $x\in\mathbb{R}^{n}$, which is easily seen to be tight when $x_{1}=\dots=x_{n}=1$. The crucial difference between this classical bound and our new inequality is the square-root improvement in the leading coefficient from $\binom{n}{\ell}$ to $\sqrt{\binom{n}{\ell}}$, which may be viewed as a consequence of the cancellations induced by the condition $\sum_{i=1}^{n}x_{i}=0$. Our new Maclaurin-type inequality, combined with the doubly centered expansion, gives rise to the first bound in Theorem 1.2.
@@ 1.3 Related work
He also developed analogous bounds for a Bernoulli version of the problem.
@@ 3.1 A warm-up example
By contrast, the above approach works directly with the ratio $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\Psi(x,\theta):=\tanh(\mu x)\frac{\theta}{\mu}$. This representation has two benefits with respect to an expansion in Hermite polynomials. First, the product structure of $\mathbb{Q}_{n}$ guarantees that for any $\theta_{i}\in\{\pm\mu\}$, the functions $\{\Pi_{i\in S}\Psi(X_{i},\theta_{i}):S\subseteq[n]\}$ are automatically orthogonal in $L^{2}(\mathbb{Q}_{n})$, so no change of measure is required. Second, and more subtly, for any $x\in\mathbb{R}^{d}$, the function $\theta\mapsto\Psi(x,\theta)$ is automatically mean-zero when $\theta\sim\mathrm{Unif}(\{\pm\mu\})$. This property guarantees that the first-order term in (3.1) vanishes, which is crucial to obtaining the correct dependence on $\mu$, and indirectly ensures that the coefficients $g_{\ell}$ are typically small. By contrast, as Section 2 makes clear, large values of the moments $m_{\alpha}$ or cumulants $\kappa_{\alpha}$ lead to the failure of the moment and cumulant methods. Further comparisons between the above two approaches are discussed in Section 5.4.

### v2

@@ 1 Introduction
Compound decisions and empirical Bayes. In the compound decision problem [Rob51], the statistician observes independent data $X_{1},\dots,X_{n}$, where $X_{i}\sim P_{\theta_{i}}$ for $i=1,\dots,n$; the goal is to estimate the vector of parameters under a separable loss $L(\theta,\widehat{\theta}):=\sum_{i=1}^{n}\ell(\theta_{i},\widehat{\theta}_{i})$. This setting is the starting point of the theory of empirical Bayes [Efr19], which views the parameters as having been drawn i.i.d. from a prior distribution, which is in turn estimated from the observations. An oracle version of this problem was suggested by Brown and Greenshtein [BG09] (see also [JZ09]), in which the statistician knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$, but not their correspondence to the observations $(X_{1},\dots,X_{n})$. Equivalently, the observations may be viewed as arising from the permutation mixture $\mathbb{P}_{n}$. As Greenshtein and Ritov [GR09] observe, however, most oracle inequalities in the empirical Bayes literature compare to an oracle that knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$ but is restricted to the use of “simple” decision rules, of the form $\hat{\theta}=(\Delta(X_{1}),\dots,\Delta(X_{n}))$ for some fixed univariate function $\Delta$. Such rules are in fact optimal when the observations arise from not from $\mathbb{P}_{n}$ but from its independent counterpart $\mathbb{Q}_{n}$. The empirical Bayes literature contains several ad hoc techniques for bounding the difference between these two approaches, going back to [HR55]. Our results can be used to give new, general risk bounds in this setting, see Corollary 1.7.
@@
Information geometry of high-dimensional mixture models. High-dimensional mixtures—such as the measures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ defined above—are ubiquitous models in theoretical statistics (see, e.g., [LNS99, CL11, IS12, WY16, JHW18]). The analysis of such measures is routine in the univariate case (or, more generally, for product measures), but obtaining sharp bounds is difficult for high-dimensional mixtures whose coordinates are not independent. The setting we study is one in which we can rigorously compare $\mathbb{P}_{n}$ to a simpler, independent counterpart $\mathbb{Q}_{n}$. This comparison may be directly relevant in analyzing mixtures with permutation invariance. Moreover, as we highlight in Section 1.2, our proofs are based on new techniques for bounding divergences between high-dimensional mixtures, which may be of broader applicability.
@@
    If $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$ and ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})<\infty$, then
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1;$ (5)
@@
Another direct corollary of Theorem 1.2 is the application of the continuity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ to mutual information with permutation structures. Let $(\theta_{1},\dots,\theta_{n})$ be a (deterministic or random) vector, with $P_{\theta_{1}},\dots,P_{\theta_{n}}\in{\mathcal{P}}$ being probability distributions over a space ${\mathcal{X}}$. Let $\vartheta=(\vartheta_{1},\dots,\vartheta_{n})$ be a uniformly random permutation of $(\theta_{1},\dots,\theta_{n})$, i.e. $\vartheta_{i}=\theta_{\pi(i)}$ with $\pi\sim\text{\rm Unif}(S_{n})$, and $X=(X_{1},\dots,X_{n})$ be a random vector with independent coordinates $X_{i}\sim P_{\vartheta_{i}}$ conditioned on $\vartheta$. From the statistical perspective, the random vector $\vartheta$ represents a permutation prior on the model parameter and $I(\vartheta;X)$ denotes the amount of information $X$ provides for estimating $\vartheta$. From the information theoretic perspective, when the vector $\theta$ represents the random message to be communicated, the maximum mutual information $I(\theta;X)$ over all possible distributions of $\theta$ is the channel capacity of the so-called noisy permutation channel [Mak20]. Our contiguity relation implies that the mutual information in both scenarios is close to its counterpart when $\vartheta\sim\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{\theta_{i}}\right)^{\otimes n}$ is i.i.d.
@@
The following inequalities hold for the mutual information $I(\vartheta;X)$ and $I(\theta;X)$:
  1. 1.
    For deterministic $\theta=(\theta_{1},\dots,\theta_{n})$,
@@
    $\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right\}.$ (7)
  2. 2.
    For possibly random $\theta=(\theta_{1},\dots,\theta_{n})$,
    $I(\theta;X)\leq\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})),$ (8)
    where $\mathrm{conv}({\mathcal{P}})$ denotes the convex hull of ${\mathcal{P}}$, and
    $N_{\mathrm{KL}}({\mathcal{P}},\varepsilon):=\min\left\{m:\inf_{{\mathcal{P}}_{0}\subseteq{\mathcal{P}}:|{\mathcal{P}}_{0}|=m}\sup_{P\in{\mathcal{P}}}\min_{P_{0}\in{\mathcal{P}}_{0}}{\sf KL}(P\|P_{0})\leq\varepsilon^{2}\right\}$
    is the $\varepsilon$-covering number of the distribution class ${\mathcal{P}}$ under the KL divergence.
In words, (1) implies that the difference in the mutual information $I(\vartheta;X)$ under a permutation prior and an i.i.d. prior on $\vartheta$ is typically independent of $n$, i.e., these priors are “equally informative” for the output $X$ in the asymptotic sense. This implication turns out to find statistical applications in problems with permutation priors; we refer to Section 5.2 for such an example in the Gaussian sequence model. The inequality (8) is an upper bound on the capacity of the noisy permutation channel, where the quantity $\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)$ is an entropic upper bound on the minimax redundancy of $\mathrm{conv}({\mathcal{P}})^{\otimes n}$ [HO97, YB99], which, by the capacity-redundancy theorem [Kem74], is the channel capacity of an i.i.d. channel $X\sim Q_{\theta}^{\otimes n}$ with $Q_{\theta}\in\mathrm{conv}({\mathcal{P}})$. We note that this entropic upper bound is tight when the KL metric entropy $\log N_{\mathrm{KL}}$ is of the same order as the Hellinger metric entropy [PW24, Theorem 32.5], which holds for Gaussian or Gaussian mixture models [Ney22, Mou23, JPW23]. The inequality (8) claims that the capacity of a noisy permutation channel is essentially upper bounded by its counterpart when the conditional distribution $P_{X|\theta}$ is i.i.d., up to an $O_{\mathcal{P}}(1)$ additive factor. This generalizes the capacity upper bound proved in [TP23] for the finite-dimensional case $\dim({\mathcal{P}})=d$, where $\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)\sim\frac{d}{2}\log n$.
@@
$\chi^{2}\left(P_{Y^{k}}\|Q_{Y^{k}}\right)\leq\frac{k(k-1)}{n(n-1)}\cdot\begin{cases}10\sum_{\ell=2}^{n}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell}&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})<\infty,\\ (e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}),\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty.\end{cases}$
In the absence of a smoothing kernel, some well-known results due to Stam [Sta78], Diaconis–Freedman [DF80], and several recent works [GK21, JGK24] yield bounds of the form ${\sf KL}(P_{X^{k}}\|Q_{X^{k}})=O(\frac{k^{2}}{n})$, which can be improved to $O(\frac{k^{2}}{n(n+1-k)})$ when $|{\mathcal{X}}|=O(1)$. In particular, $P_{X^{k}}$ and $Q_{X^{k}}$ are, in general, statistically indistinguishable when $k=o(\sqrt{n})$, and that this requirement can be improved to $k=o(n)$ when ${\mathcal{X}}$ is small. [^1] Note that if no restriction is made on the size of ${\mathcal{X}}$, then the birthday paradox shows that the requirement $k=o(\sqrt{n})$ is tight. In comparison, Theorem 1.5 implies that as long as the exchangeable distribution can be written as a convolution with a kernel ${\mathsf{K}}$ satisfying $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$ (such as a Gaussian), the mixtures are statistically indistinguishable for any $k=o(n)$, without any restriction on $|{\mathcal{X}}|$. This observation suggests that the addition of noise in the form of the kernel ${\mathsf{K}}$ has a similar qualitative effect as a bound on the size of ${\mathcal{X}}$; we leave investigation of this interesting connection to future work. A similar upper bound on the KL divergence for the noisy setting also appeared in [TP23, Theorem 2], with a linear dependence on both $\frac{k}{n}$ and the cardinality of ${\mathcal{P}}$; we will provide a detailed comparison in Section 1.3.
@@
$\left(\mathbb{E}_{\mathbb{P}_{n}}[f]-\mathbb{E}_{\mathbb{Q}_{n}}[f]\right)^{2}\leq\mathbb{E}_{\overline{P}^{\otimes n}}\left[f^{2}\right]\cdot\frac{12{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n}.$
Theorem 1.6 shows that the squared TV distance between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of the order $O_{\mathcal{P}}(\frac{1}{n})$. In addition, the mean difference of any test function $f$ under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of order $O_{\mathcal{P}}(n^{-1/2})$ as long as the second moment of $f$ under the i.i.d. distribution $\overline{P}^{\otimes n}$ is bounded.
@@
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq M\sqrt{6n{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}.$
Corollary 1.7 shows that for a large class of models and losses, asymptotically these two types of decision rules achieve similar mean risks $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(\sqrt{n})$, strengthening a “folklore” result $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=o(n)$ in [HR55]. For the quadratic loss $L(\vartheta,\widehat{\vartheta})=\|\vartheta-\widehat{\vartheta}\|_{2}^{2}$, the risk difference can be improved to $O(1)$ via a different strategy, as shown for a class of models in [GR09]; we will discuss this improvement in Section 5.3.
@@ 1.2 Proof techniques
$|e_{\ell}(x)|\leq C\sqrt{\binom{n}{\ell}}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}\right)^{\ell/2}.$ (10)
This bound may be contrasted with Maclaurin’s inequality [Mac30], a version of which implies the bound $|e_{\ell}(x)|\leq\binom{n}{\ell}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}\right)^{\ell/2}$ for all $x\in\mathbb{R}^{n}$, which is easily seen to be tight when $x_{1}=\dots=x_{n}=1$. The crucial difference between this classical bound and our new inequality is the square-root improvement in the leading coefficient from $\binom{n}{\ell}$ to $\sqrt{\binom{n}{\ell}}$, which may be viewed as a consequence of the cancellations induced by the condition $\sum_{i=1}^{n}x_{i}=0$. Our new Maclaurin-type inequality, combined with the doubly centered expansion, gives rise to the first bound in Theorem 1.2.
@@ 1.3 Related work
He also developed analogous bounds for a Bernoulli version of the problem. A dimension-independent upper bound on the KL divergence ${\sf KL}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ was also observed and obtained in [TP23, Theorem 2] for a finite class ${\mathcal{P}}$, used as a crucial technical step towards establishing the capacity upper bound of the noisy permutation channel in [Mak20]. Their proof relies on an anti-concentration property for the histogram of independent discrete observations with non-identical distributions, leading to an upper bound linear in $|{\mathcal{P}}|$. They also prove a bound analogous to the one presented in Theorem 1.5 for $k$-dimensional marginals with $k<n$ via a convexity argument [TP23, Appendix A]; however, their bound scales as $O(\frac{k}{n})$ rather than $O(\frac{k^{2}}{n^{2}})$.
@@ 3.1 A warm-up example
By contrast, the above approach works directly with the ratio $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\Psi(x,\theta):=\tanh(\mu x)\frac{\theta}{\mu}$. This representation has two benefits with respect to an expansion in Hermite polynomials. First, the product structure of $\mathbb{Q}_{n}$ guarantees that for any $\theta_{i}\in\{\pm\mu\}$, the functions $\{\Pi_{i\in S}\Psi(X_{i},\theta_{i}):S\subseteq[n]\}$ are automatically orthogonal in $L^{2}(\mathbb{Q}_{n})$, so no change of measure is required. Second, and more subtly, for any $x\in\mathbb{R}^{d}$, the function $\theta\mapsto\Psi(x,\theta)$ is automatically mean-zero when $\theta\sim\mathrm{Unif}(\{\pm\mu\})$. This property guarantees that the first-order term in (3.1) vanishes, which is crucial to obtaining the correct dependence on $\mu$, and indirectly ensures that the coefficients $g_{\ell}$ are typically small. On the other hand, as Section 2 makes clear, large values of the moments $m_{\alpha}$ or cumulants $\kappa_{\alpha}$ lead to the failure of the moment and cumulant methods. Further comparisons between the above two approaches are discussed in Section 5.4.